A circular lake is located at a certain barangay in Tagaytay. An existing road runs tangential to the circular lake. Twelve kilometers from the point of tangency, along the existing road, a new road of 8 km length is constructed to the periphery of the lake at B. The new road is extended to cross the lake. Compute the length of the bridge needed to cross the lake
Answer #1

First we sketch the situation as done below

P 8km B 12km e ? е С. ΑΙ o

Given that the sand existing soad AP is tangential to the circular lake. Length AP = 12 km From point p a road (new) of 8 km

Now observe that <ABOTSOBC = (90-0) +0 = 90° = (ABC CABP = 180 - <ABC A ABP = 180-90°-90° right angled triangle. Thus, ✓ 1445

The answer is Bridge length is 10km.

Kindly give a thumbs up. This is nothing but appreciation that you are benefiting.